A firm produces 500 units of output. At this production level the firm’s marginal cost of production is £180 and the firm’s total cost of production is £55,500. At this level of production, the firm exhibits increasing or decreasing returns to scale, show your working


1
Expert's answer
2022-01-17T16:22:54-0500

Average total cost is ATC = TC/Q = 55,500/500 = 111.

As MC > ATC, then the firm exhibits increasing returns to scale, because ATC starts increasing after ATC = MC, which is upward-sloping.